OpenStudy (anonymous):

Find the following for the two functions: -domain and range -x and y intercepts -horizontal asymptote(s) -vertical asymptote(s) 1. 2/x^2-2x-3 2. x^2+x-2/x^2-3x-4

OpenStudy (anonymous):

So far I think I am correct, but idk 1. domain: 2 +/- sqrt3/2 range: all real numbers x intercept: none y intercept: -.67 horizontal: 0 Vertical: -1

OpenStudy (anonymous):

\[\frac{2}{(x-3)(x+1)}\] domain is all numbers except \(x=3,x=-1\)

OpenStudy (anonymous):

no \(x\) intercept because it is never zero (since the numerator is 2)

OpenStudy (anonymous):

oh okay and I think that there is another vertical intercept, but I cannot seem to find it

OpenStudy (anonymous):

if \(x=0\) then \(y=-\frac{2}{3}\) for the \(y\) intercept

OpenStudy (anonymous):

vertical asymptotes at the zeros of the denominator, namely \(x=-1,x=3\)

OpenStudy (anonymous):

horizontal asymptote what you said, \(y=0\)

OpenStudy (anonymous):

alright for the secong equation is it domain and range is all real numbers

OpenStudy (anonymous):

range is not all real numbers however

OpenStudy (anonymous):

for example , this can never be 0, so zero is not in the range i believe the range is \((-\infty,-\frac{1}{2})\cup (0,\infty)\) but that takes some checking

OpenStudy (anonymous):

easy with calc, somewhat harder with algebra (i am still talking about the first one)

OpenStudy (anonymous):

\[y=\frac{2}{x^2-2x+3}\] \[y(x^2-2x-3)=2\] \[yx^2-2yx-(3y+2)=0\] solve the quadratic for \(x\)

OpenStudy (anonymous):

okay for the equation 2 \[\frac{ x^2+x-2 }{ x^2-3x-4 }\] would the domain be all real numbers?

OpenStudy (anonymous):

no, the denominator cannot be zero

OpenStudy (anonymous):

\(x^2-3x-4=0\) or \((x-4)(x+1)=0\) so domain does not contain \(4\) or \(-1\)

OpenStudy (anonymous):

oh okay is y intercept: 1/2 x intercept: -1.5 and 3.5

OpenStudy (anonymous):

it is probably easier if you write this in factored form as \[\frac{(x+2)(x-1)}{(x-4)(x+1)}\] then you can find what you need

OpenStudy (anonymous):

two \(x\) intercepts, at \(x=-2,x=1\)

OpenStudy (anonymous):

Horixontal asymptote is 1 and vertical is -1 and 0 right?